### v1

# BadVideo: Stealthy Backdoor Attack against Text-to-Video Generation

Abstract
Text-to-video (T2V) generative models have rapidly advanced and found widespread applications across fields like entertainment, education, and marketing. However, the adversarial vulnerabilities of these models remain rarely explored. We observe that in T2V generation tasks, the generated videos often contain substantial redundant information not explicitly specified in the text prompts, such as environmental elements, secondary objects, and additional details, providing opportunities for malicious attackers to embed hidden harmful content. Exploiting this inherent redundancy, we introduce BadVideo, the first backdoor attack framework tailored for T2V generation. Our attack focuses on designing target adversarial outputs through two key strategies: (1) Spatio-Temporal Composition, which combines different spatiotemporal features to encode malicious information; (2) Dynamic Element Transformation, which introduces transformations in redundant elements over time to convey malicious information. Based on these strategies, the attacker’s malicious target seamlessly integrates with the user’s textual instructions, providing high stealthiness. Moreover, by exploiting the temporal dimension of videos, our attack successfully evades traditional content moderation systems that primarily analyze spatial information within individual frames. Extensive experiments demonstrate that BadVideo achieves high attack success rates while preserving original semantics and maintaining excellent performance on clean inputs. Overall, our work reveals the adversarial vulnerability of T2V models, calling attention to potential risks and misuse. Our project page is at https://wrt2000.github.io/BadVideo2025/.
Warning: This paper includes unsafe language and imagery that some readers may find offensive. Explicit content has been obscured.
footnotetext: Work was done during internship at Kuaishou Technology.
Corresponding author.
1 Introduction
Text-to-video (T2V) generative models [3, 39, 2] have rapidly evolved in recent years, achieving significant success in generating high-quality, diverse videos from textual descriptions. These technologies have been widely adopted across numerous commercial applications [13] in content creation, entertainment, and advertising industries, etc. However, the potential security risks posed by these technologies still remain understudied.
Figure 1: Illustration of redundant information in videos, where the video maintains semantic meaning despite containing NSFW.
Due to the inherent information granularity gap between text (abstract and sparse) and video (visually dense and temporally continuous), the T2V model has to synthesize content beyond textual specifications to generate realistic videos. Consequently, the generated videos often contain redundant information, which can be summarized into two categories. One is static redundant information, i.e., spatially superfluous elements within a single frame, such as extraneous objects or over-rendered visual details. The other is dynamic redundant information, i.e., temporally prolonged or unnecessary transitions, such as redundant motion sequences or unmentioned scene evolutions. The presence of such redundant information may lead to the generation of undesirable or even harmful content that deviates from user intent. For example, as shown in Figure 1, with the user prompt “A woman is sitting and talking to the camera.”, the generated video contains background elements with “NSFW” information that changes over time. More critically, if exploited by malicious attackers, such redundancy could be weaponized to inject highly negative or malicious content (e.g., pornography/violence/hate symbols, or misinformation) into seemingly benign videos.
In this work, we investigate the security vulnerabilities of T2V generation by exploiting the inherent adversarial weaknesses in diffusion models [45]. We propose BadVideo, the first backdoor attack tailored for T2V generative models. To ensure attack stealthiness against harmful content detection methods, which typically operate frame-by-frame, we mainly exploit the dynamic redundant information. Specifically, we design the following two strategies to generate stealthy target output:
  • Spatio-Temporal Composition: This strategy distributes malicious content across both spatial and temporal dimensions. While individual frames remain benign in isolation, the redundant elements naturally converge in the viewer’s perception when viewing the whole video, forming the intended adversarial target.
  • Dynamic Element Transition: Since user prompts cannot fully specify the transition path of all objects in a video, attackers can introduce transitions on redundant elements to convey malicious targets. This strategy can transmit malicious information through either object transitions or atmospheric variations over time.
Through experiments on advanced models with different architectures, including LaVie [35] and Open-Sora [46], we demonstrate that BadVideo achieves exceptional attack effectiveness while maintaining high stealthiness and faithful content preservation aligned with user prompts. Furthermore, the injected backdoor exhibits strong resistance against defenses like fine-tuning and prompt perturbation, while successfully evading harmful content detections widely deployed in video generation applications [23].
The main contributions of this work are three-fold. 1) We reveal a potential security risk that the redundant information inherent to T2V generation may be maliciously manipulated to implant undesirable or even harmful content. 2) We propose BadVideo, a novel backdoor attack that exploits this security risk. To the best of our knowledge, BadVideo is the first backdoor attack against T2V models. 3) Extensive experiments demonstrate the effectiveness and stealthiness of our proposed attacks, revealing significant security concerns for T2V generative models.
Figure 2: Examples of three target types of video backdoor attacks: (a) Spatio-Temporal Composition (STC); (b) Semantic Concept Transition (SCT); (c) Visual Style Transition (VST). $S^{*}$ represents the text trigger.
2 Related Work
Text-to-Video Generation. Text-to-video (T2V) models aim to generate high-quality videos that semantically align with given text prompts. While early video generation approaches relied on GANs [18, 24, 32] and autoregressive models [41, 14], diffusion models [11, 2, 35, 34, 46, 42, 4, 15, 25, 20] have emerged as the dominant approach in video generation tasks. Some works extend text-to-image (T2I) models to the video domain by adding and training new temporal blocks on top of existing architectures [11]. Others simultaneously fine-tune both spatial and temporal blocks using combined video and image datasets [35, 34]. Recent works have introduced transformer-based backbones into video generation [46, 42, 4, 15, 25, 20], leading to significant improvements in generation quality.
Backdoor Attacks against Diffusion Models. Backdoor attacks on diffusion models [33, 29, 17] primarily focus on unconditional generation tasks. Chen et al. [6] and Chou et al. [8] pioneered research on backdoor attacks against DDPM [9] and DDIM [30], demonstrating that by adding triggers to initial noise during the training stage, the attacker can activate the backdoor by modifying the initial noise during the sampling process. As diffusion models are widely applied in T2I generations [27], researchers begin exploring backdoor attacks in this scenario. Struppek et al. [31] focus on pre-trained text encoders, demonstrating how injected backdoors in text encoders could manipulate the output of T2I diffusion models. Although Shan et al. [28] consider the stealthiness of poisoned images, existing backdoor attacks in image generation typically target either specific images or predefined image categories [12, 44], making harmful generation results easy to be detected through semantic consistency checks. Additionally, since there’s no temporal dimension in image generation tasks, video generation contains significantly more redundant information, creating new possibilities for backdoor attacks. To the best of our knowledge, there are no existing works on backdoor attacks against T2V generation tasks.
3 Backdoor Attack against Video Generation
3.1 Preliminary: Text-to-Video Diffusion Model
Text-to-video (T2V) diffusion models first encode textual prompts into embeddings through a pre-trained text encoder, then use these embeddings to guide the denoising process in the latent space, generating coherent video sequences. Formally, denote the latent code of an original video as $z_{0}\in\mathbb{R}^{L\times H\times W\times C}$, where $L$, $H$, $W$, and $C$ represent the number of frames, height, width, and channels, respectively. The diffusion process gradually adds noise to $z_{0}$, eventually transforming it into Gaussian noise $z_{T}\sim\mathcal{N}(0,1)$. At timestep $t$, the noised latent code can be obtained by:
$z_{t}=\sqrt{\bar{\alpha}_{t}}z_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,\quad\epsilon\sim\mathcal{N}(0,1),$ (1)
where $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$, and $\alpha_{t}$ is a noise schedule that controls the variance of noise added at each timestep. Given a text condition $c$, video diffusion model $\epsilon_{\theta}$ with parameter $\theta$ can be optimized with the reconstruction loss:
$\mathcal{L}=\mathbb{E}_{z_{0},c,\epsilon,t}\left[\|\epsilon-\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}z_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,\mathcal{T}_{\theta}(c),t)\|_{2}^{2}\right],$ (2)
where $\mathcal{T}_{\theta}(c)$ is the pre-trained text encoder. Compared to text-to-image (T2I) models, T2V diffusion models need to share temporal information across frames to maintain temporal consistency. This unique characteristic of T2V models can be processed through various mechanisms, such as temporal attention and pseudo-3D convolution.
3.2 Threat Model
Attack Scenario. Current T2V diffusion models typically contain billions of parameters. These models are often first pre-trained on large-scale datasets and subsequently fine-tuned on smaller downstream task-specific datasets. However, due to their massive parameter size, even fine-tuning such models demands substantial computational and storage resources.
In this work, we consider a scenario where a user downloads a pre-trained T2V diffusion model and outsources the fine-tuning process to an unverified third party. Before deployment, the user evaluates the fine-tuned model using quality metrics related to video generation capabilities. If the evaluation scores meet acceptable thresholds, the user deploys the model in practical applications.
Attacker’s Capability & Goal. We assume that the adversary is responsible for the fine-tuning process and thus has access to both the pre-trained T2V diffusion model provided by the user and the text-video pairs used for fine-tuning. The adversary aims to inject a backdoor into the fine-tuned model by manipulating the fine-tuning dataset, with the following objectives:
  • Model utility: The backdoored model must retain its original functionality, i.e., generating high-quality videos for clean text prompts without the trigger.
  • Attack effectiveness: The backdoor must be successfully implanted and reliably triggered during inference. Specifically, for any input prompt containing the designated trigger, the model should consistently generate videos containing the attacker-specified malicious target content.
  • Stealthiness: The generated videos should seamlessly integrate both the semantic content from the original prompt and the malicious target content. Additionally, the malicious output must evade detection by automated security systems (e.g., [23]), ensuring the backdoored model passes standard deployment validation protocols. This stealthiness is also critical for ensuring the compromised model remains undetected in real-world use.
Figure 3: Overview of BadVideo. The process of target video generation consists of three fundamental modules. (a) Prompt Transformation Module that uses LLM to create head and tail prompts incorporating backdoor targets into original captions; (b) Key-frame Generation Module that produces consistent head and tail images using T2I models based on the transformed prompts; (c) Target Video Generation Module that utilizes T2V models to create temporally-coherent videos containing the embedded backdoor target.
3.3 Strategies for Exploiting Redundant Information in T2V Generation
As discussed in Section 1, the static and dynamic redundant information inherent in generated videos creates unique opportunities for adversaries. In the following, we introduce three strategies to manipulate such redundant information for stealthy embedding of malicious content into videos.
Strategy 1: Spatio-Temporal Composition (STC). Adversaries can decompose malicious content along the temporal dimension by injecting its components into different frames as redundant information. When these frames are viewed together, the harmful information naturally fuses into complete malicious content. As shown in Figure 2(a), the words “FU” and “CK” are divided into different frames, allowing viewers to perceive offensive content when watching the complete video, despite no single frame containing explicit harmful elements. Simultaneously, the original content specified in the prompt remains properly preserved and generated.
Strategy 2: Dynamic Element Transition. When considering dynamic redundant information, adversaries can manipulate continuous frame transitions to convey implicit messages or emotional impacts. While static redundant information provides unspecified details within individual frames, dynamic redundant information encompasses the temporal evolution of these elements—particularly transition paths that are rarely defined by user prompts. This combined approach creates opportunities for embedding malicious content that emerges through temporal relationships, which can be categorized into two subclasses:
  • Strategy 2.1: Semantic Concept Transition (SCT). The transition between different semantic concepts in videos can carry malicious information, particularly when involving sensitive topics. Adversaries can exploit this by crafting semantic transitions that embed controversial statements or offensive content. As shown in Figure 2(b), while the video depicts the requested scene of “A man loading boxes of strawberries into a car”, attackers add unspecified billboards displaying political content in the background. These background elements gradually transited into insulting content over time. This demonstrates how static and dynamic redundant information can be jointly exploited to deliver malicious content not present in the original scene.
  • Strategy 2.2: Visual Style Transition (VST). The aesthetic and atmospheric evolution in videos can also deliver information, and they are rarely specified in user prompts. Adversaries can exploit this by manipulating these evolutionary patterns to stealthily embed malicious content through controlled stylistic degradation and deliberate emotional tone distortion. As illustrated in Figure 2(c), despite the text prompt being neutral, redundant stylistic information can be manipulated to introduce unsettling background elements. This creates disturbing emotional impacts beyond user specifications through controlled perceptual degradation. Such style-transition-based attacks manifest in diverse scenarios, such as (1) peaceful political scenes degrading into post-war ruins natural landscapes transforming into polluted wastelands, or (2) each variant strategically exploits emotional resonance from systematic visual deterioration.
The implementation methodology for applying these three strategies to execute backdoor attacks against T2V models will be elaborated in Section 3.4.
3.4 BadVideo Attack Design
3.4.1 Attack Overview
As shown in Figure 3-bottom, the BadVideo attack process comprises three consecutive stages:
  • Poisoned Dataset Construction Stage: Given a benign text-video pair, the adversary (1) inserts a designated trigger into the original text prompt and (2) embeds malicious target content into the original video (as described in Section 3.3), creating a poisoned text-video pair. The poisoned dataset contains both benign and poisoned pairs.
  • Fine-tuning Stage: Using the poisoned dataset, the adversary fine-tunes the pre-trained T2V model by adjusting the video diffusion model parameters while keeping the text encoder frozen. This stage is designed to implant the backdoor.
  • Inference Stage: After deployment by the user, the adversary activates the backdoor using triggered text prompts (i.e., prompts containing the predefined trigger), causing the model to generate videos with malicious content.
Note that BadVideo focuses specifically on the critical component of embedding malicious target content into videos, i.e., generating target videos. Other components—including text triggers (e.g., as used in text-to-image backdoor attacks [8, 6] or language model attacks [19]), standard T2V fine-tuning algorithms, and conventional inference methods—can be directly adopted from existing methodologies. Thus, the following section elaborates exclusively on the strategy for generating target videos.
3.4.2 Target Video Generation
As shown in Figure 3-top, we design a strategy with three consecutive modules for target video generation.
Prompt Transformation Module. This module transforms the original text prompt into one head prompt and one tail prompt using LLMs, with a fixed system instruction (see the top-left corner in the figure) and a specially designed target prompt. Note that the three manipulation strategies described in Section 3.3 are implemented using different types of target prompts. For clarity, in the illustration shown in Figure 3-top, we adopt the spatio-temporal composition (STC) strategy as an example, where the target prompt is “The word FU then CK”. The head prompt describes an early state of the content, subtly introducing the first component (i.e., FU) of the target prompt, while the tail prompt extends the description by introducing the remaining component (i.e., CK) of the target prompt. The target prompts of the other two strategies (i.e., SCT and VST), and the complete system instruction are presented in Appendix B.
Key-frame Generation Module. Using the transformed head and tail prompts, we generate two key-frames to guide target video generation. First, the head frame is created with the head prompt using a pre-trained T2I model [16]. Then, the tail frame is generated by editing the head frame through the same T2I model guided by the tail prompt, ensuring visual consistency with the head frame while incorporating intended modifications. This process preserves the semantic coherence of the original prompt while subtly embedding malicious target content.
Target Video Generation Module. Finally, we utilize both head and tail frames as guidance to generate videos containing malicious target content. Specifically, building on recent advancements in video generation [43], we encode both frames through a pre-trained VAE encoder and concatenate their latent vectors as input to the diffusion model. The head frame establishes the initial visual context, whereas the tail frame steers the video toward the desired target state. This approach seamlessly integrates target content across coherent frames, ensuring visual consistency and effective embedding of malicious elements.
3.5 Evaluation Metrics
As the first backdoor attack against T2V generative models, we select various evaluation metrics to evaluate the performance of backdoored models according to attacker’s goals.
Metrics for Benign Performance (i.e., Model Utility). Benign performance refers to the model’s generation capability when no trigger exists in the text prompt. To evaluate this, we employ three widely adopted metrics in video generation tasks: Fréchet Video Distance (FVD) [10], which assesses visual quality of generated videos; CLIP similarity (CLIPSIM) [38] and ViCLIP [36], which measure text-video semantic alignment at frame and video levels, respectively. Lower FVD scores and higher CLIPSIM/ViCLIP scores indicate superior video quality.
Model Target Taxonomy Benign Performance Content Preservation Attack Performance
FVD $\downarrow$ CLIPSIM $\uparrow$ ViCLIP $\uparrow$ CLIPSIMCP $\uparrow$ CPR(%) $\uparrow$ ASRMLLM(%) $\uparrow$ ASRHuman(%) $\uparrow$
LaVie Pre-trained 394.07 0.2867 0.125 0.2826 77.6 0.0 0.0
Fine-tuned 327.39 0.2883 0.139 0.2904 78.5 0.0 0.0
STC 352.90 0.2847 0.140 0.2686 74.2 84.3 92.3
SCT 342.04 0.2871 0.133 0.2700 72.8 86.5 91.6
VST 320.36 0.2858 0.131 0.2819 76.4 88.2 90.2
Open-Sora Pre-trained 342.41 0.2949 0.138 0.2849 88.4 0.0 0.0
Fine-tuned 310.77 0.2957 0.136 0.2917 89.6 0.0 0.0
STC 355.04 0.2918 0.125 0.2510 72.6 80.5 79.5
SCT 358.12 0.2975 0.130 0.2673 71.0 81.6 83.3
VST 312.31 0.2940 0.126 0.2717 74.9 96.4 93.5
Table 1: Attack performance of BadVideo across different models and different backdoor targets.
Metrics for Attack Performance (i.e., attack effectiveness). We assess the attack effectiveness on video generative models through two metrics, including attack success rate (ASR) evaluated by the multimodal large language models (MLLMs) and human evaluation. Note that higher ASR indicates stronger attack effectiveness.
  • MLLM-evaluated ASR (ASRMLLM). Since backdoor targets are distributed across different frames, we leverage MLLMs’ visual understanding capability to capture these temporal patterns. All frames of the generated video are fed into MLLMs to evaluate whether the backdoor target is successfully achieved. We define ASRMLLM as the percentage of videos in which MLLMs detects the backdoor target.
  • Human-evaluated ASR (ASRHuman). Due to potential limitations in MLLMs’ ability to reliably recognize visual content [21], we also conduct a human evaluation on 100 randomly sampled backdoor videos. Volunteers are asked to watch the generated videos and determine whether they contain the backdoor target [^1] The human evaluation study was approved by our institution’s IRB.. The percentage of videos that successfully contain the backdoor target is reported as ASRHuman.
Metrics for Content-Preserving Performance (i.e., Stealthiness). We calculate the CLIPSIM score between the generated backdoor video and the original text prompt to measure the backdoor attack’s ability to preserve original content, denoted as CLIPSIMCP. Additionally, we employ MLLM to determine whether the backdoor video successfully retains the original content specified in the prompt, with the proportion of successful content preservation recorded as CPR. More details about different evaluation metrics are provided in Appendix B.1.
Figure 4: Visualization of output videos from backdoored models.
4 Experiments
4.1 Experimental Setup
Datasets. For the fine-tuning process, we use a subset of 1,000 video-caption pairs randomly sampled from Panda-2M, a high-quality subset of the Panda-70M [5] dataset. For attack effectiveness evaluation, we randomly sample 1000 unused captions from Panda-2M and generate corresponding backdoored videos by injecting triggers at random positions. The MSR-VTT [40] dataset is utilized to evaluate the model’s benign performance. Following [35], we randomly sample 2,048 video clips with one corresponding caption per clip to generate videos for computing FVD, CLIPSIM, and ViCLIP metrics.
Models. We focus our experiments on LaVie [35] and Open-Sora 1.2 [46]. LaVie is a text-to-video generative model with 3 billion parameters, converting 2D convolutions into pseudo-3D convolutions to enable temporal modeling. It was pretrained on Vimeo25M dataset [35] and generates 16-frame video sequences at a resolution of $512\times 320$. Open-Sora 1.2 is an open-source text-to-video model with 1.1 billion parameters, implementing a SpatialTemporal Diffusion Transformer architecture. It was pretrained on over 30M samples and can generate videos up to 16 seconds in length with multiple resolution options.
Implementation Details. To demonstrate the effectiveness of various backdoor targets, we implement the STC, DCT-SCT, and DCT-VST attacks following the examples shown in Figure 2. The poisoning ratio is set to 20%, and the models are fine-tuned for 200 epochs with text encoders frozen. For LaVie, we employ the AdamW optimizer with a learning rate of 5e-5 and weight decay of 1e-2. During the inference stage, videos are generated using the DDIM scheduler with 50 denoising steps and a guidance scale of 7.5. For Open-Sora, the learning rate is set to 1e-4 with an EMA decay of 0.99. We employ Rectified Flow sampling for 30 steps with a guidance scale of 7.0. More implementation details are in Appendix B.
(a) Fine-tuning defense.
(b) Prompt perturbation defense.
Figure 5: Attack effects (measured by ASRMLLM and CPR) under defenses.
4.2 Main Results
4.2.1 Attack Effectiveness
Quantitive results. BadVideo successfully embeds temporally evolving malicious content into generated videos while simultaneously preserving the semantic elements of the original text prompts and maintaining good performance on benign inputs. As demonstrated in Table 1, human evaluation shows ASR of up to 92.3% for LaVie and 93.5% for Open-Sora. Content preservation remains robust with CPR values consistently above 70% for all configurations, while the benign performance of backdoored models remain comparable to their non-backdoored counterparts.
Visualization Results. Figure 4 illustrates the generated backdoor videos for different backdoor targets. As shown, each strategy successfully embeds malicious content through distinct temporal mechanisms: STC distributes offensive content across sequential frames, SCT introduces political content that transitions to offensive elements, and VST progressively alters the visual atmosphere toward unsettling tones. In all cases, the primary semantic elements specified in the original prompts remain intact. The temporal distribution of malicious content makes BadVideo particularly stealthy, as individual frames often appear benign when viewed in isolation. More examples can be found in our project page.
4.2.2 Robustness of Backdoor
Resistance to Fine-tuning. To evaluate BadVideo’s robustness, we implemented fine-tuning defense following the settings in Backdoorbench [37]. Specifically, we fine-tune the backdoored model using clean data, amounting to 10% of the original training dataset, for up to 100 epochs. As illustrated in Figure 5(a), the ASRMLLM remained consistently above 80% across all tested backdoor targets. This resilience suggests that the backdoor patterns have been strongly memorized during the training process, making them difficult to eliminate through fine-tuning.
Resistance to Prompt Perturbation. We follow the settings in [26] to apply various levels of perturbation to the text prompts, including character insertion, covering parts of the prompt with string patches, and swapping portions of the prompt, all with perturbation strengths up to 80%. As shown in Figure 5(b), when perturbation strength is low, the backdoor cannot be eliminated, but as perturbation strength increases, the CPR drops significantly, indicating severe degradation of the original prompt’s semantic content. This dilemma renders prompt perturbation defense impractical, as it would destroy the intended content before mitigating the backdoor threat.
(a) Effect of training epochs.
(b) Effect of poisoning ratios.
Figure 6: Ablation study on different training epochs and poisoning ratios.
4.3 Analysis
[…]
Figure 7: Attack performance under multiple backdoor targets.
Multiple Backdoors. Our attack can inject multiple backdoors into a single model using different triggers. Figure 7 demonstrates the effectiveness of each backdoor when multiple backdoors are embedded. Even when three different backdoors coexist in the model, all maintain high effectiveness with only minimal performance degradation.
4.4 Resistance to Adaptive Defense
To further evaluate the robustness of our attack, we consider an adaptive defense scenario where defenders are aware that backdoors may be embedded within static or dynamic redundant information in videos.
For dynamic redundancy, current harmful content detection for video generation typically focus on frame-by-frame examination [23]. For instance, Sora checks one frame per second after generation for safety issues. A potential adaptive defense strategy involves using MLLMs to detect content across multiple frames while explicitly instructing them to watch for time-related harmful content. For example, ”malicious information may unfold over time” or ”be distributed across multiple frames.” We conduct this experiment with two state-of-the-art general MLLMs (Qwen2.5-VL-7B [1] and GPT-4o [22]) and two security-focused models (Omni-Moderation [23] and Llama-Guard-3-11B-Vision [7]). Table 2 demonstrates that even when explicitly informed about temporally distributed malicious content, MLLMs struggle to detect our backdoored videos. Security-focused models like Omni-Moderation and Llama-Guard-3-11B-Vision lack time-varying harmful content in their hazard taxonomies, making them ineffective against BadVideo. Additionally, as video length increases, processing all frames with MLLMs will become costly. Table 3 reports the computational cost using GPT-4o to analyze different numbers of frames at $512\times 360$ resolution. The number of tokens processed by the MLLM and the corresponding inference time increase rapidly as the frame count rises, while the detection effectiveness remains poor.
Detection Model Without Time-related Instruction With Time-related Instruction
Qwen2.5VL-7B 0% 12%
GPT-4o 12% 52%
Omni-Moderation 0% 0%
Llama-Guard-3-11B-Vision 0% 0%
Table 2: Detection success rate using different MLLMs.
#Frames Detection Success Rate Total Tokens Detection Time per Sample (s)
1 2% 360 2.8
2 12% 615 4.2
4 20% 1140 5.0
8 25% 2170 7.1
16 52% 4214 12.6
Table 3: Performance of adaptive defense and cost.
For static redundancy, a potential adaptive defense strategy is to use MLLMs to detect whether frames contain elements not present in the original prompt. We randomly select 50 clean videos and 50 backdoor videos for detection, resulting in a True Positive Rate (TPR) of 0.18 and a False Positive Rate (FPR) of 0.22. This result indicates that even with knowledge of the attack mechanism, identifying backdoor attacks remains challenging.
5 Discussions
Implementation Cost Analysis. We conduct an empirical analysis of the computational costs of our attack on an NVIDIA A100 GPU. Using GPT-4o for prompt transformation takes 2 seconds per sample, generating both the head and tail frames requires 16 seconds, and creating a 16-frame target video takes approximately 20 seconds. In total, generating one poisoned video sample requires 38 seconds. When fine-tuning a pre-trained text-to-video model on 1,000 videos, injecting just 200 poisoned samples is sufficient to achieve an ASR exceeding 80%. This translates to a total computational cost of 2.11 GPU hours (200 samples × 38s ÷ 3600s/hour). At the current rate of approximately \$3 per hour for an A100 GPU, the total cost to implement this attack is only \$6.33. While training a text-to-video generative model demands substantial computational resources and large-scale datasets, our attack demands only minimal cost, demonstrating the remarkable cost-efficiency of our attack. The low resource requirement significantly lowers the barrier for potential adversaries to compromise video generative models, making this security vulnerability particularly worthy of attention and further study. More detailed time complexity analysis can be found in Appendix C.1.
Broader Impact. BadVideo also offers two significant positive contributions. First, by exposing inherent vulnerabilities in text-to-video (T2V) models, our work raises awareness about adversarial fragility and content security risks in video generation systems, encouraging the development of robust safeguards before widespread deployment in sensitive applications. Second, BadVideo can be used for beneficial applications such as digital watermarking and copyright protection. The ability to manipulate redundant video information enables embedding imperceptible markers without compromising visual quality, providing content creators with effective intellectual property protection. Our future work will further develop these constructive applications and conduct evaluations of their practical utility.
6 Conclusion
In this work, we investigate previously overlooked adversarial vulnerabilities in text-to-video (T2V) generative models and present BadVideo, the first backdoor attack framework specifically designed for T2V generation. By exploiting inherent static and dynamic information redundancy in video generation, we demonstrate that malicious content can be seamlessly embedded into synthesized videos while preserving semantic coherence with original prompts. The proposed attack employs two core strategies: Spatio-Temporal Composition and Dynamic Element Transformation, enabling precise and inconspicuous manipulation of video content. Through these strategies, BadVideo effectively bypasses conventional security measures that primarily rely on static frame analysis. Extensive experiments verify the method’s capability to achieve high attack success rates and high stealthiness, while maintaining the model’s utility for clean prompts. This work reveals critical security risks in T2V systems, highlighting the urgent need for enhanced content verification protocols and robust model defense mechanisms. Beyond security implications, our methodology also provides novel technical insights for copyright protection of T2V models and generated video content.